DETERMINATION OF THE STRUCTURE OF CYCLOPENTENEOXIDE AND THE ARGON CYCLOPEN-TENEOXIDE VAN DER WAALS COMPLEX

\underline{ANDREA J. MINEI}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; JENNIFER VAN WIJNGAARDEN, Department of Chemistry, Mount Holyoke College, South Hadley, MA 01075; WALLACE C. PRINGLE, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

Rotational spectra of cyclopenteneoxide and the argon cyclopenteneoxide van der Waals complex were studied using pulsed-jet Fabry-Perot Fourier transform microwave spectroscopy. Spectra of the normal isotopomer along with those of the ^{13}C singly substituted isotopomers for both the monomer and the complex were measured in the frequency region of 6-20 GHz. The boat structure for cyclopenteneoxide was confirmed with naturally abundant ^{13}C isotopic substitution. For the argon cyclopenteneoxide complex, both a and b-type transitions were observed and the rotational constants for the all ^{12}C isotopomer were determined to be A = 3268.2537(5), B = 993.3458(2), and C = 950.4300(2) MHz. The coordinates of the argon in the principal axis system of cyclopenteneoxide are a = 0.27, b = 0.42, and c = 3.91 Å, such that the argon atom is exo to the bridged oxygen of the ring. An investigation into whether or not the argon atom undergoes large amplitude cross-ring motion will be presented.